For the following fusion reaction, calculate the c
Answer #1

Δm = m(2-H) + m(2-H) - m(3-H) + m(1-H)
= 2.01410 + 2.01410 - 3.01605 - 1.00782
= 0.00433 u
= 0.00433 g/mol
= 4.33*10^-6 kg/mol

ΔE = Δm*c^2
=(4.33*10^-6)*(3*10^8)^2
= 3.897*10^11 J/mol

Answer: 3.897*10^11 J/mol